OpenStudy (anonymous):

how can i solve 4^{3x=1}=16,384?

OpenStudy (amistre64):

using properties of logs and or exponents

OpenStudy (anonymous):

exponent

OpenStudy (amistre64):

\[\Large B^{m+n}=B^m*B^n\]

OpenStudy (amistre64):

divide of the 4^(=1?) part and try to get the right side into a base 4 equivalent

OpenStudy (amistre64):

another thought is that since 4 = 2^2 we can try to turn it all into base 2 stuff

OpenStudy (anonymous):

\[4^{3x+1}=16,384\]

OpenStudy (amistre64):

use the property i started with\[4^{3x+1}=4^{3x}*4^1\]

OpenStudy (amistre64):

divide off the 4^1 from each side to lessen that number

OpenStudy (amistre64):

4021 ---------- 4 ) 16384

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

\[4^{3x}=4021\] is what we have left to figure on

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

if you are just forced to deal with exponents; lets write up a base 4 table ^5 ^4 ^3 ^2 ^1 ^0 1024 256 64 16 4 1

OpenStudy (amistre64):

4^6 = 4096

OpenStudy (anonymous):

i see

OpenStudy (amistre64):

and it helps if i learn to divide; 16384/4 = 4096 too

OpenStudy (amistre64):

so, that leaves us with\[4^{3x}=4^6\] equate exponents 3x=6 when x=?

OpenStudy (anonymous):

so x= 6?

OpenStudy (amistre64):

3x = 6

OpenStudy (anonymous):

ok thx